OpenStudy (anonymous):

The measure of a complement of an angle is three more than twice the measure of the angle. Find the measures of the angle and the complement. help me find a formula! :D

OpenStudy (anonymous):

angle complement is 90-angle. so: (90-x)-2x=3

OpenStudy (anonymous):

90-x-2x=3 3x=87 x=87/3=29 complement =(90-29)=61

OpenStudy (anonymous):

wait, how did you get (90-x)-2x=3, I understand that (90-x) is ONE of the complements, but the other complement is 2x+3 right? how did you get -2, and the = 3?

OpenStudy (anonymous):

hello..?